Circular No. 6831

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998S IN NGC 3877 A. Nakamura, Kuma, Ehime, Japan, reports that a prediscovery image

of SN 1998S was obtained by K. Kawakami, Kochi University, on Mar. 2.70 UT with a 0.20-m f/6.3 reflector + unfiltered CCD; Nakamura derived a magnitude of V ∼ 16.0 (difficult to measure because the star lies on the bright rim of the host galaxy). D. D. Balam, University of Victoria, reports the following precise position for SN 1998S from an R-band CCD image: α = 11^{h}46^{m}06^{s}.18, δ = +47^{o}28′55′′.5 (equinox 2000.0), which is 13′′.6 west and 46′′ south of the host galaxy’s center; a nearby star (R = 16.4) has position end figures 16^{s}.96, 25′36′′.8. J.-C. Merlin, Le Creusot, France, reports the following position end figures for SN 1998S, obtained with a 0.40-m f/5 reflector + CCD: 05^{s}.85, 54′′.9 (unfiltered mag 13.6 on Mar. 5.844).

SUPERNOVA 1987A IN THE LARGE MAGELLANIC CLOUD D. Baade, European Southern Observatory (ESO); and L. Wang, D. A.

Howell, and J. C. Wheeler, University of Texas, report: “SN 1987A was observed on Mar. 2 and 3 with the ESO 3.6-m telescope (+ EFOSC2). A low-resolution spectrum (range 429.9–817.6 nm; resolving power ∼ 1000; p.a. of the slit 220^{o}; width of the slit 1′′) shows dramatic changes of the line profiles compared to those obtained a year earlier. The most remarkable features are: 1) The Hα profile is now much broader; the blue wing of the Hα line has grown much stronger than a year ago and has now merged with the neighboring [O I] line at 630.0 and 636.4 nm, and is stronger than the same [O I] line; the red wing of Hα is also much stronger and broader, being traceable at least to 14 000 km/s. 2) Many new narrow emission lines were detected, which have a spatial extension comparable to that of the inner bright ring. The blobs with the highest S/N ratios appear near Hβ in wavelength and are not seen at Hα. They are unresolved at our spectral resolution. At least six distinct narrow features were detected on top of the broad Hβ emission line of the supernova ejecta. Comparison with earlier data suggests that the interaction between the SN 1987A ejecta and its circumstellar material is increasing in strength rapidly. Observations in all accessible wavelength ranges are strongly encouraged.”

V4334 SAGITTARII Visual magnitudes estimates via J. Mattei, AAVSO (cf. IAUC 6825):

Feb. 24.1 UT, 12.6 (B. Monard, Pretoria, S. Africa); 25.73, 12.8 (R. Stubbings, Drouin, Australia); Mar. 1.74, 12.8 (Stubbings); 3.07, 12.8 (Monard).

1998 March 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green